OpenStudy (anonymous):

what is the steps and process to find the point slope form from 4,4 to 6,8

OpenStudy (jdoe0001):

http://mathb.in/12751

OpenStudy (jdoe0001):

hmm the point slope... form... ok... that only gets you the slope so

OpenStudy (jdoe0001):

http://mathb.in/12752

OpenStudy (anonymous):

@texaschic101 do you know it

OpenStudy (texaschic101):

first we need to find the slope by using the slope formula. slope = (y2 - y1) / (x2 - x1) (4,4) x1 = 4 and y1= 4 (6,8) x2 = 6 and y2 = 8 now we sub slope = (8 - 4) / (6 - 4) slope = 4/4 which equals 1 now we use the point-slope formula : y - y1 = m(x - x1) slope(m) = 1 you can pick either set of points... (4,4) x1 = 4 and y1 = 4 now we sub y - 4 = 1(x - 4) ==> this is your point slope form or this point... (6,8) x1 = 6 and y1 = 8 slope(m) = 1 y - y1 = m(x - x1) y - 8 = 1(x - 6) ==> this is also point-slope form It does not matter which points you pick, the ending answer when solved will result in the same line. any questions ? oh, and also, putting it in point-slope form does not mean solving it. Just put it in that form, parenthesis and all.

OpenStudy (texaschic101):

you also got to be careful if you have negative numbers in your sets of points...either when finding the slope or putting it in point-slope form. Sometimes you will come across something like this... 2 - (-3) or -3 - (-2). Just pay attention to your signs, keeping in mind that two like signs when multiplied together are positive.

OpenStudy (anonymous):

:)))))) you are awesome thank youuuuu

OpenStudy (texaschic101):

glad to help :)